OpenStudy (anonymous):

Use the distribute property to simplify the expression. 4(2x-4) Hay can some one help me on this?

OpenStudy (anonymous):

8 • (x - 2)

OpenStudy (anonymous):

It would be 8x-16

OpenStudy (anonymous):

Hmm.

OpenStudy (anonymous):

Simplify : 4 • (2x - 4) Step 1 : Multiply 4 by 2x-4 Pulling out like terms : 1.1 Pull out like factors: 2x - 4 = 2•(x - 2) Final result : 8 • (x - 2)

OpenStudy (anonymous):

think your wrong Dwight

OpenStudy (anonymous):

Let me check again . Because DWIGHT is NEVER WRONG.

OpenStudy (anonymous):

I'm telling you it's 8x-16

OpenStudy (mathstudent55):

Here's the distributive property of multiplication over subtraction: \(a(b - c) = ab - ac\) Here's an example: Use the distributive porperty to simplify: \(3(5x - 9) \) \(= 3 \cdot 5x - 3 \cdot 9 \) \(= 15x - 27 \) Now study the rule and the example and try to apply it to your problem.

OpenStudy (anonymous):

Nevermind. It is 8x-16. I was wrong

OpenStudy (anonymous):

Think your right ur answer danirod16 its one of the options

OpenStudy (anonymous):

It is that, I promise

OpenStudy (anonymous):

Uhm , yes we all agree it is 8x-16. You welcome guys no need to thank me. Just doing my job as a Schrute.

OpenStudy (anonymous):

Okay. Thanks allot your help is greatly appreciated. ^_^

OpenStudy (anonymous):

Don't thank me.. Thank my father.

OpenStudy (anonymous):

ill post another if im in dire help thx allot. and don't worry were ganna have allot of time together while waiting for the money

OpenStudy (anonymous):

Waiting for the money? Are you speaking of my Schrute bucks?

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

Well if you were wondering we Schrutes use every part of the goose. The meat has a delicious smoky rich flavor. Plus, you can use the molten goose grease and save it in the refrigerator, thus saving you a trip to the store for a can of expensive goose grease.

OpenStudy (anonymous):

you thought way up there in the sky that you were safe. but dont here amigo you hit the ground

OpenStudy (anonymous):

Heres a Schrute buck. and if You're quoting Far cry know that video games corrupt the brain..

OpenStudy (anonymous):

LOOK me in the eye. LOOk ME IN the F**King eye

OpenStudy (anonymous):

Vaas from Far Cry 3 and Dwight from the office having a conversation? time to screen-shot!

OpenStudy (anonymous):

STFU STFU STFU D:

OpenStudy (anonymous):

That just happened :l

OpenStudy (anonymous):

Come any closer and im going to free spur soul.

OpenStudy (anonymous):

DO YOU WANT ME TO SLICE YOU OPEN LIKE I DID YOUR FRIEND

OpenStudy (anonymous):

Noo :3

OpenStudy (anonymous):

As a farmer I know that when an animal is sick sometimes the right thing to do is put it out of it’s misery. With the electricity we are using to keep Vaas alive in Far cry we could power a small fan for two days. You tell me what’s unethical